The length of a rectangular garden is 8 ft less than 4 times its width. If the perimeter of the garden is 52 ft, find the dimensions of the garden.
Answer by checkley75(3666)

LENGTH=4W-8
WIDTH=W
PERIMETER=2*L+2*W
52=2(4W-8)+2W
52=8W-16+2W
10W=52+16
10W=68
W=68/10
W=6.8 ANSWER FOR THE WIDTH.
L=(4*6.8-8)
L=27.2-8
L=19.2 ANSWER FOR THE LENGTYH.
PROOF
2*19.2+2*6.8=52
38.4+13.6=52
52=52